C–H...π and C–F...H–C INTERACTIONS IN THE ACETYLENE–FLUOROFORM DIMER

\underline{SEAN A. PEEBLES}, MICHAL M. SERAFIN, REBECCA A. PEEBLES, Department of Chemistry, Eastern Illinois University, 600 Lincoln Ave., Charleston, IL 61920 USA.

Rotational spectra for four isotopomers of the HCCH–HCF_{3} complex have been identified by Fourier-transform microwave spectroscopy. The spectra exhibit considerable fine splittings, presumably arising from the internal motion of one or both subunits; both first- and second-order Stark effects have been observed. A fit of selected K=0 and 1 components for the normal isotopomer (assigned using second-order Stark effect data) gave an rms of around 4 kHz and provided rotational constants A = 5450.236(4) MHz, B = 1381.3361(9) MHz, C = 1375.1333(12) MHz. These rotational constants and the planar moments are consistent with a structure of C_{s} symmetry exhibiting C–H...π and C–F...H–C interactions and are in excellent agreement with a structure obtained from optimizations at the MP2/6-311++G(2d,2p) level.